OpenStudy (drummmlers):

Hana Juarez drove for 80 miles in heavy traffic. She then reached the highway where she drove 260 miles at an average speed that was 25 miles per hour greater than the average speed she drove in heavy traffic. If the total trip took 6 hours, determine her average speed in heavy traffic and on the highway.

OpenStudy (raffle_snaffle):

Do you have any ideas?

OpenStudy (faiqraees):

Distance at heavy traffic/Speed +Distance at highway/Speed = Total time \[\large\rm \frac{ 80}{x }+\frac{260}{x+25} = 6\] Now solve the quadratic equation